in a certain region of space, the electric potential is V(x,y)= Axy-By2+Cy whereA,B,C are positive constants with values 5,4,10 units. a) calculate the x,y,z componet of eletric field for x= 2 V/m and y= 1V/m. b) also find the resultant electirc field
Answer #1

a) Ex = -dV/dx

= -d(Axy-By^2 + Cy) /dx

= -Ay = -4y

Ey = -dV/dy

= -d(Axy-By^2 + Cy)/dy

= -Ax - 2By + C

= -5x - 2*4*y + 10

= -5x - 8y + 10

Ez = -dV/dz

= -d( Axy - By^2+ Cy) / dz.

= 0

b) electric field = -4y i + ( -5x - 8y + 10) j